# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v2

@@ V.1 Reconstructed differential SMFs and UVLFs for each present-day classification
The proto-MW/LG SMFs/UVLFs have shapes similar to the universal SMF/UVLF from FIREboxHR, lending confidence to the assumption that the proto-MW/LG provides a representative sample of the global average SMF/UVLF during reionization. However, the amplitude of the proto-MW/LG SMF/UVLF is lower than that of FIREboxHR, particularly in the isolated suite. This could be attributed to the difference in the calculation of the volume that goes into the denominator while computing the SMFs/UVLFs (see section V). The stellar masses and UV magnitudes in FIREboxHR are also computed differently than in this work (see Feldmann et al. 2025 for details). Additionally, the FIREboxHR box has a higher mean matter density (exactly matching the mean density of the Universe) across the simulation volume than the MW-like zoomed simulations at high redshifts, which could also lead to this difference in amplitude. This discrepancy is smaller in the case of the paired LG-like suite because these simulations have a higher mean matter density than the isolated MW-like simulations, and the galaxies form stars at earlier times in the paired suite, leading to more star-forming galaxies at high redshifts. We plan to explore this difference further in future work.
@@
Figure 7: Differential SMFs at $z=7$ in the fossil record scenario. Similar to Figure 5, but for the SMFs in the fossil record scenario (curves with shading). This scenario assumes all the stellar mass from multiple progenitors in enclosed in a single one, as we would infer observationally. The SMFs and ratios for the all progenitors scenario shown in Figure 5 are included for comparison (solid curves with no shading). Unlike the all progenitors scenario, the fossil record scenario has a mass-dependent bias: the reconstruction is less complete for lower-mass galaxies than for higher-mass galaxies. The effect is most prominent in sample III (green), followed by samples II (blue) and I (magenta) respectively.
@@
Figure 8: Differential UVLFs at $z=7$ in the fossil record scenario. Similar to Figure 7, but for the differential UVLFs. The mass-dependence of the reconstruction noted in Figure 7 for the SMFs is less pronounced for the UVLFs, especially in the LG-like systems.
@@
The overall shapes of the reconstructed SMFs/UVLFs from the three samples are qualitatively similar to that of the proto-MW/LG. Adding the progenitors of the stellar streams recovers the SMF/UVLF of the proto-MW/LG more accurately, particularly in terms of normalization, up to $\sim 50$%. Adding the progenitors of the phase-mixed galaxies improves the reconstruction even more. Including the disrupted systems helps most in recovering intermediate- and high-mass progenitors that are missed when using only intact low-mass galaxies. Furthermore, the halo-to-halo scatter in samples II and III generally appears lower than sample I, with the exception of bottom right panels in Figures 5 and 6. But there the scatter is from just 3 paired halos and hence may not be robust. This is even more apparent when the low-mass/faint-end slope is estimated (Section VI.1). Adding more disrupted galaxies to our present-day samples leads to more robust recovery of our proto-MW/LG SMFs/UVLFs by potentially reducing the halo-to-halo scatter, which is sensitive to the variations in accretion histories. This is particularly prominent in the case of the isolated suite, since it probes a wider range of accretion histories than the paired suite, leading to more scatter in the case of Sample I.
@@
The ratios of the reconstructions from each sample to the proto MW/LG are nearly flat over a wide range of stellar masses/magnitudes. The scatter is larger in the high-mass end because of low-number statistics. Samples I and II have a slight bias at lower (fainter) masses (magnitudes), but remain almost flat at the low-mass/faint-end ($M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ or $M_{\mathrm{UV}}\gtrsim-14$). Sample II recovers $\sim$50% of the proto-MW/LG SMF/UVLF while sample I recovers only $\sim$25%, a factor of $\sim$2 improvement in the normalization. Sample III does even better, recovering $\sim$75% of the proto-MW/LG with reduced bias. These results argue that the progenitors of the stellar streams and phase-mixed galaxies are crucial for a more accurate and precise representation of the proto-MW/LG and the universal SMFs/UVLFs at high redshifts.
@@
V.3 Reconstructed differential SMFs and UVLFs from the fossil record of present-day substructures
@@
We next reconstruct the SMFs from the fossil record, by combining the stellar masses of all the progenitors of each galaxy into a single progenitor (Figure 7). We use this combined stellar mass to calculate UV magnitudes and reconstruct the UVLFs (Figure 8). The color scheme in these figures remains the same as in Figure 5, but the dashed lines now show the median SMFs/UVLFs from the fossil record, and the shaded regions show the 68% halo-to-halo scatter in the fossil record. We show the SMFs/UVLFs from the all progenitors scenario as solid lines for reference but (for the sake of readability) do not show their halo-to-halo scatter.
@@
Using the fossil record instead of the complete merger history has several effects on reconstructing the $z=7$ SMFs/UVLFs. First, the overall normalization is reduced because there are fewer progenitors in the fossil record scenario. Second, the SMFs/UVLFs from the fossil record scenario in Figures 7 and 8 (top panel) and their ratios with respect to that of the proto-MW/LG (bottom panel) decrease towards the low-mass/faint-end because many of the low-mass/faint progenitors are combined into a single more massive progenitor in the fossil record assumption. Since the present-day disrupted galaxies can be more massive than intact galaxies, and thus can have multiple progenitors at high redshifts, the fossil record approach leads to larger differences in shape towards the low-mass/faint-end when they are included. Furthermore, the number of massive systems with multiple high-$z$ progenitors also increases when we add more types of present-day substructures. Hence the SMFs/UVLFs of sample III, which contains the most substructures in general as well as the largest number of disrupted galaxies, show the largest difference, followed by sample II and I, respectively.
@@
Except for the case of the sample III SMF in the isolated suite, the ratios in the fossil record scenario remain nearly flat between $10^{5}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ for the other SMFs. In other words, the bias is not mass dependent over this range. The UVLFs between $14\lesssim M_{\mathrm{UV}}\lesssim-10$ also remain nearly flat. However, over the entire mass/UV magnitude range, the ratios decrease towards the low-mass/faint-end. This occurs because the fossil record scenario combines multiple progenitors into a single effective progenitor, which reduces the number of low-mass/faint-end progenitors. The difference is exacerbated for sample II and sample III which have better completeness at higher stellar mass. We also note that the relative shape of the SMFs/UVLFs from the three samples do not change significantly in the case of the proto-LG, particularly for the UVLFs. These results do not rule out the possibility of using disrupted galaxies along with intact galaxies in the near-far technique, but rather indicate that our inferences are sensitive to the selection criteria for both the present-day substructures and their high-$z$ progenitors, along with the effects of the environment they reside in. They are also sensitive to whether we use the SMF or the UVLF. We explore this selection-dependence more quantitatively in the next section.
@@
Figure 9: Adding streams and phase-mixed galaxies improves the accuracy and precision of the recovered slope of the differential SMF at low masses. We compare the low-mass slopes of the differential SMFs at $z=7$ for samples I (magenta), II (blue), and III (green) to both the full SMF of the proto-MW/LG (black) and the universal slope estimate from FIREboxHR (yellow). Errorbars indicate the standard deviation on the recovered slope for each individual halo, while the horizontal lines show the median ($\tilde{\mu}_{\alpha}$) of the bootstrapped sample of mean $\alpha$ values across all the halos, computed as described in Section VI.1 and Appendix C. The shaded regions show the median ($\tilde{\sigma}_{\alpha}$) of the bootstrapped sample of standard deviations in $\alpha$, representing the halo-to-halo scatter. We show ($\tilde{\mu}_{\alpha}$) and ($\tilde{\sigma}_{\alpha}$) in the rightmost points for proto-MW and proto-LG by combining them as well as separately. In the all progenitors scenario (top), sample III consistently reproduces the low-mass-end slope of the proto-MW/LG SMF most accurately ($\alpha$ closest to that of the proto-MW/LG for each individual halo) and precisely (smaller errorbars for each individual halo). The median slope ($\tilde{\mu}_{\alpha}$) for sample III is also closest to that of the proto-MW/LG and have the smallest halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$), followed by sample II and I. In the fossil record scenario (bottom), samples II and III still have consistently smaller uncertainties for each individual halo but sample I is more accurate than II and III at recovering the slope in some of the halos. The conclusions remain similar when MW and LG are separately considered.
@@ VI Shapes of the SMFs and UVLFs during the EoR
VI.1 Low-mass/faint-end slopes
@@
Since a primary goal of the near-far technique is to determine the low-mass end slope of the SMF at $z\gtrsim$ 6, we next consider the effect of including the progenitors of the disrupted galaxies on the recovery of this slope. Gandhi et al. (2024) showed that the low-mass-end slope can be accurately recovered from the fossil record progenitors of just the intact low-mass galaxies, without taking disrupted galaxies into account. They compute the low-mass-end slope by fitting a single power-law to the SMF across a mass range of $_{\star}=10^{4.5-6.5} MM_{\odot}$ and report the median slope across all of the simulations, concluding that the slope inferred from the fossil record of just the intact low-mass galaxies agrees with that of the proto-MW/LG. We test this result by calculating the slopes of the reconstructed high-$z$ SMFs/UVLFs for our three samples and the proto-MW/LG, but fit a Schechter function to incorporate massive galaxies as well (Schechter, 1976). This is well-motivated by observations (Duncan et al., 2014; Bouwens et al., 2015; Finkelstein et al., 2015; Song et al., 2016; Atek et al., 2018; Ishigaki et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2021, 2023; Whitler et al., 2025), but alternative models (e.g., double power-law) also exist, particularly for explaining the excess number of galaxies at the bright-end ($\mathrm{M_{UV}\lesssim-22})$ (Bowler et al., 2015, 2017, 2020; Donnan et al., 2023; Harikane et al., 2023; Sun et al., 2023b; Harikane et al., 2024; Donnan et al., 2024; Chemerynska et al., 2025), which is beyond the brightest set of galaxies in our samples. Instead of fitting the Schechter functions to binned SMFs/UVLFs, we perform Bayesian modeling using the PyMC package (Abril-Pla et al., 2023) to estimate the posterior distribution of the slope parameter $\alpha$. This approach eliminates the need for binning, which can affect the result for $\alpha$, and provides a full posterior distribution for $\alpha$. In the case of FIREboxHR we do use the binned SMFs/UVLFs, but generate posteriors of $\alpha$ by fitting Schechter functions to these SMFs/UVLFs using the same approach. The details of our methodology, along with illustrative results, are described in Appendix B.
@@
We estimate the posterior distribution of the slope $\alpha$, which follows an approximately Gaussian distribution (see Appendix B), and calculate its mean and standard deviation for the three samples in each simulation. Figures 9 and 10 show the slopes obtained by our fits for the SMFs and UVLFs respectively, compared to the ones from the proto-MW/LG and the universal SMFs/UVLFs at $z=7$. The all progenitors scenario is shown in the top panel, while the fossil record scenario is in the bottom panel. The points and the error bars in each figure denote the mean and standard deviation of the posterior distribution for the low-mass/faint-end slope $\alpha$ for the proto-MW/LG (black), sample I (red), sample II (blue) and sample III (green) for each simulation. The mean slope of the universal SMF/UVLF is shown by the yellow line, while the shaded region shows the standard deviation. To estimate population-level parameters, we bootstrap the individual $\alpha$ values, resampling from their Gaussian posteriors. This yields: (1) a central slope estimate ($\tilde{\mu}_{\alpha}$) - median of the bootstrapped means; (2) halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$) - median of the bootstrapped standard deviation. We show these by the horizontal lines and the shaded regions, respectively, in Figures 9 and 10. To compare the results between proto-MW and proto-LG, we separately do the analysis for the isolated and paired galaxy suites. For clarity, we also show these estimates in the rightmost points of both the figures. We emphasize that $\tilde{\sigma}_{\alpha}$ is not the standard deviation of the bootstrapped mean $\alpha$ distribution, but rather an average range of possible $\alpha$ values across all the halos in the respective galaxy suite, defining the halo-to-halo scatter. We discuss the details of our bootstrapping method in Appendix C. Tables 2 and 3 summarizes our slope estimates. We note that the median values for the isolated and paired galaxy suites calculated separately, in general, provide similar conclusions as the median values for the 10 halos combined. However, the results for the paired galaxy suite are derived from only three simulations and thus may not be robust. Since the trend is similar we discuss our results for $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$ when all the halos are combined, unless stated otherwise.
@@
The low-mass end slope values of the proto-MW/LG and sample I reported here are different than the ones in Gandhi et al. (2024) by $\approx$0.1, likely due to the difference in the slope estimation method. In addition, some of the present-day intact galaxies in Gandhi et al. (2024) are instead identified as part of our set of disrupted galaxies, so our Sample I is not identical to the set used in that paper.
@@
We find several key features when we include the disrupted galaxy progenitors while computing the low-mass-/faint-end slopes. First, when we compare the individual low-mass/faint-end slopes from the SMFs/UVLFs of the three samples with the one from the proto-MW/LG, we find that in most halos the slope of the proto-MW/LG SMF/UVLF is consistently better reproduced by sample III, followed by sample II and I, respectively. This is also reflected in the median values, $\tilde{\mu}_{\alpha}$, from our bootstrapped samples. Second, including more disrupted galaxies reduces the standard deviation of $\alpha$ values from sample I within each halo, by up to $\sim$30% in sample II and $\sim$50% in sample III. This is because the increased number of progenitors improves statistical robustness. Third, there is a scatter in the values of the slope across different halos, which implies that different accretion histories can lead to different values of the reconstructed low-mass-/faint-end slope, leading to poorly constrained values across the realizations. However, we find that this halo-to-halo scatter can be reduced by adding the disrupted galaxies, as shown by the decrease in the median $\tilde{\sigma}_{\alpha}$ values (rightmost points). This reduction in the scatter, in comparison to sample I, can be $\sim$20% in case of sample II and $\sim$40% in case of sample III. The effect of an unusual accretion history in producing a poorly constrained slope can thus be mitigated by adding the disrupted galaxies.
@@
We reach similar conclusions when we consider the effect of adding disrupted galaxies separately to a MW-like environment and a LG-like environment. The halo-to-halo scatter in case of a LG-like environment is smaller compared to that of a MW-like environment, but the results are derived from only three simulations and may not be robust. Additionally, the low-mass-end for sample I and III is steeper than proto-LG because of the R&R paired halo (see Figure 9). Furthermore, the slopes are susceptible to how present-day and high-$z$ galaxies are selected, which we discuss in VI.2 and hence the should be treated with caution.
@@
Interestingly, the low-mass-end slopes of the SMFs (see Figure 9) and the faint-end slopes of the UVLFs (see Figure 10) differ. The UVLFs have a shallower slope at the faint-end ($\approx$ -1.4) compared to the low-mass-end of the SMFs ($\approx$ -1.7). Even though extrapolations of observed UVLFs predict a faint-end slope of $\approx$ -2, results from FIREboxHR also seem to be pointing to a shallower slope (see Table 2). Although this could depend on the specific distribution used to draw the progenitor masses and the slope estimation method, the difference between the SMF and UVLF slopes invites further investigation. Using the power-law nature of the SMF at the low-mass-end and the mass-to-light ratios computed from our progenitor stellar halos, we find that the shallower slope in the UVLFs, can be partially explained due to the superlinear dependence of UV luminosity with stellar mass. Since young, massive O/B stars generate most of the UV radiation and lower-mass galaxies are less efficient at producing them, this leads to suppressed UVLFs at the faint-end. We refer the reader to Appendix D for further details and figures demonstrating our analysis. In any case, we better reproduce the proto-MW/LG SMFs/UVLFs after including present-day disrupted galaxies.
@@
Another important difference can be discerned between the all progenitors scenario and the fossil record scenario: the latter case results in flatter slopes - by $\sim$0.07 on average across the three samples when both the isolated and paired galaxy suites are combined. As discussed in Section V.3, this effect is due to combining multiple progenitors into a single one in the fossil record scenario and it gets amplified for samples II and III as they have more massive substructures with multiple progenitors. The level of flattening varies depending on how both the present-day and high-$z$ progenitor galaxies are selected.
@@
Sample Stellar Mass Functions UV Luminosity Functions
@@
Sample Stellar Mass Functions UV Luminosity Functions

### v3

@@ V Stellar Mass Functions and Ultraviolet Luminosity Functions at High-Redshifts
The proto-MW/LG SMFs/UVLFs have shapes similar to the universal SMF/UVLF from FIREboxHR, lending confidence to the assumption that the proto-MW/LG provides a representative sample of the global-average SMF/UVLF during reionization. However, the amplitude of the proto-MW/LG SMF/UVLF is lower than that of FIREboxHR, particularly in the isolated suite. This could be attributed to the difference in the calculation of the volume that goes into the denominator while computing the SMFs/UVLFs (see section V). The stellar masses and UV magnitudes in FIREboxHR are also computed differently than in this work (see Feldmann et al. 2025 for details). Additionally, the FIREboxHR box has a higher mean matter density (exactly matching the mean density of the Universe) across the simulation volume than the MW-like zoomed simulations at high redshifts, which could also lead to this difference in amplitude. This discrepancy is smaller in the case of the paired LG-like suite because these simulations have a higher mean matter density than the isolated MW-like simulations, and the galaxies form stars at earlier times in the paired suite, leading to more star-forming galaxies at high redshifts. We plan to explore this difference further in future work.
@@
Figure 7: Differential SMFs at $z=7$ in the fossil-record scenario. Similar to Figure 5, but for the SMFs in the fossil-record scenario (curves with shading). This scenario assumes all the stellar mass from multiple progenitors in enclosed in a single one, as we would infer observationally. The SMFs and ratios for the all-progenitors scenario shown in Figure 5 are included for comparison (solid curves with no shading). Unlike the all-progenitors scenario, the fossil-record scenario has a mass-dependent bias: the reconstruction is less complete for lower-mass galaxies than for higher-mass galaxies. The effect is most prominent in Sample III (green), followed by Samples II (blue) and I (magenta) respectively.
@@
Figure 8: Differential UVLFs at $z=7$ in the fossil-record scenario. Similar to Figure 7, but for the differential UVLFs. The mass-dependence of the reconstruction noted in Figure 7 for the SMFs is less pronounced for the UVLFs, especially in the LG-like systems.
@@
The overall shapes of the reconstructed SMFs/UVLFs from the three samples are qualitatively similar to that of the proto-MW/LG. Adding the progenitors of the stellar streams recovers the SMF/UVLF of the proto-MW/LG more accurately, particularly in terms of normalization, up to $\sim 50$%. Adding the progenitors of the phase-mixed galaxies improves the reconstruction even more. Including the disrupted systems helps most in recovering intermediate- and high-mass progenitors that are missed when using only intact low-mass galaxies. Furthermore, the halo-to-halo scatter in Samples II and III generally appears lower than sample I, with the exception of bottom right panels in Figures 5 and 6. But there the scatter is from just 3 paired halos and hence may not be robust. This is even more apparent when the slopes at the low-mass/faint end is estimated (Section VI.1). Adding more disrupted galaxies to our present-day samples leads to more robust recovery of our proto-MW/LG SMFs/UVLFs by potentially reducing the halo-to-halo scatter, which is sensitive to the variations in accretion histories. This is particularly prominent in the case of the isolated suite, since it probes a wider range of accretion histories than the paired suite, leading to more scatter in the case of Sample I.
@@
The ratios of the reconstructions from each sample to the proto MW/LG are nearly flat over a wide range of stellar masses/magnitudes. The scatter is larger in the high-mass end because of low-number statistics. Samples I and II have a slight bias at lower (fainter) masses (magnitudes), but remain almost flat at the low-mass/faint end ($M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ or $M_{\mathrm{UV}}\gtrsim-14$). Sample II recovers $\sim$50% of the proto-MW/LG SMF/UVLF while Sample I recovers only $\sim$25%, a factor of $\sim$2 improvement in the normalization. Sample III does even better, recovering $\sim$75% of the proto-MW/LG with reduced bias. These results argue that the progenitors of the stellar streams and phase-mixed galaxies are crucial for a more accurate and precise representation of the proto-MW/LG and the universal SMFs/UVLFs at high redshifts.
@@
V.3 Reconstructed Differential Stellar Mass Functions and Ultraviolet Luminosity Functions from the Fossil Record of Present-day Substructures
@@
We next reconstruct the SMFs from the fossil record, by combining the stellar masses of all the progenitors of each galaxy into a single progenitor (Figure 7). We use this combined stellar mass to calculate UV magnitudes and reconstruct the UVLFs (Figure 8). The color scheme in these figures remains the same as in Figure 5, but the dashed lines now show the median SMFs/UVLFs from the fossil record, and the shaded regions show the 68% halo-to-halo scatter in the fossil record. We show the SMFs/UVLFs from the all-progenitors scenario as solid lines for reference but (for the sake of readability) do not show their halo-to-halo scatter.
@@
Using the fossil record instead of the complete merger history has several effects on reconstructing the $z=7$ SMFs/UVLFs. First, the overall normalization is reduced because there are fewer progenitors in the fossil-record scenario. Second, the SMFs/UVLFs from the fossil-record scenario in Figures 7 and 8 (top panel) and their ratios with respect to that of the proto-MW/LG (bottom panel) decrease towards the low-mass/faint end because many of the low-mass/faint progenitors are combined into a single more massive progenitor in the fossil-record assumption. Since the present-day disrupted galaxies can be more massive than intact galaxies, and thus can have multiple progenitors at high redshifts, the fossil record approach leads to larger differences in shape towards the low-mass/faint end when they are included. Furthermore, the number of massive systems with multiple high-$z$ progenitors also increases when we add more types of present-day substructures. Hence the SMFs/UVLFs of Sample III, which contains the most substructures in general as well as the largest number of disrupted galaxies, show the largest difference, followed by Sample II and I, respectively.
@@
Except for the case of the Sample III SMF in the isolated suite, the ratios in the fossil-record scenario remain nearly flat between $10^{5}\,M_{\odot}\lesssim M_{\star}\lesssim 10^{6.5}\,M_{\odot}$ for the other SMFs. In other words, the bias is not mass dependent over this range. The UVLFs between $14\lesssim M_{\mathrm{UV}}\lesssim-10$ also remain nearly flat. However, over the entire mass/UV-magnitude range, the ratios decrease towards the low-mass/faint end. This occurs because the fossil-record scenario combines multiple progenitors into a single effective progenitor, which reduces the number of low-mass/faint progenitors. The difference is exacerbated for Sample II and Sample III which have better completeness at higher stellar mass. We also note that the relative shape of the SMFs/UVLFs from the three samples do not change significantly in the case of the proto-LG, particularly for the UVLFs. These results do not rule out the possibility of using disrupted galaxies along with intact galaxies in the near-far technique, but rather indicate that our inferences are sensitive to the selection criteria for both the present-day substructures and their high-$z$ progenitors, along with the effects of the environment they reside in. They are also sensitive to whether we use the SMF or the UVLF. We explore this selection-dependence more quantitatively in the next section.
@@
Figure 9: Adding streams and phase-mixed galaxies improves the accuracy and precision of the recovered slope of the differential SMF at low masses. We compare the low-mass slopes of the differential SMFs at $z=7$ for Samples I (magenta), II (blue), and III (green) to both the full SMF of the proto-MW/LG (black) and the universal slope estimate from FIREboxHR (yellow). Errorbars indicate the standard deviation on the recovered slope for each individual halo, while the horizontal lines show the median ($\tilde{\mu}_{\alpha}$) of the bootstrapped sample of mean $\alpha$ values across all the halos, computed as described in Section VI.1 and Appendix C. The shaded regions show the median ($\tilde{\sigma}_{\alpha}$) of the bootstrapped sample of standard deviations in $\alpha$, representing the halo-to-halo scatter. We show ($\tilde{\mu}_{\alpha}$) and ($\tilde{\sigma}_{\alpha}$) in the rightmost points for proto-MW and proto-LG by combining them as well as separately. In the all-progenitors scenario (top), Sample III consistently reproduces the low-mass-end slope of the proto-MW/LG SMF most accurately ($\alpha$ closest to that of the proto-MW/LG for each individual halo) and precisely (smaller errorbars for each individual halo). The median slope ($\tilde{\mu}_{\alpha}$) for Sample III is also closest to that of the proto-MW/LG and have the smallest halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$), followed by Sample II and I. In the fossil-record scenario (bottom), Samples II and III still have consistently smaller uncertainties for each individual halo but Sample I is more accurate than II and III at recovering the slope in some of the halos. The conclusions remain similar when MW and LG are separately considered.
@@ VI Shapes of the SMFs and UVLFs during the EoR
VI.1 Slopes at the Low-mass/Faint end
@@
Since a primary goal of the near-far technique is to determine the low-mass-end slope of the SMF at $z\gtrsim$ 6, we next consider the effect of including the progenitors of the disrupted galaxies on the recovery of this slope. Gandhi et al. (2024) showed that the low-mass-end slope can be accurately recovered from the fossil record progenitors of just the intact low-mass galaxies, without taking disrupted galaxies into account. They compute the low-mass-end slope by fitting a single power law to the SMF across a mass range of $_{\star}=10^{4.5-6.5} MM_{\odot}$ and report the median slope across all of the simulations, concluding that the slope inferred from the fossil record of just the intact low-mass galaxies agrees with that of the proto-MW/LG. We test this result by calculating the slopes of the reconstructed high-$z$ SMFs/UVLFs for our three samples and the proto-MW/LG, but fit a Schechter function to incorporate massive galaxies as well (Schechter, 1976). This is well-motivated by observations (Duncan et al., 2014; Bouwens et al., 2015; Finkelstein et al., 2015; Song et al., 2016; Atek et al., 2018; Ishigaki et al., 2018; Bhatawdekar et al., 2019; Bouwens et al., 2021, 2023; Whitler et al., 2025), but alternative models (e.g., double power law) also exist, particularly for explaining the excess number of galaxies at the bright end ($\mathrm{M_{UV}\lesssim-22})$ (Bowler et al., 2015, 2017, 2020; Donnan et al., 2023; Harikane et al., 2023; Sun et al., 2023b; Harikane et al., 2024; Donnan et al., 2024; Chemerynska et al., 2025), which is beyond the brightest set of galaxies in our samples. Instead of fitting the Schechter functions to binned SMFs/UVLFs, we perform Bayesian modeling using the PyMC package (Abril-Pla et al., 2023) to estimate the posterior distribution of the slope parameter $\alpha$. This approach eliminates the need for binning, which can affect the result for $\alpha$, and provides a full posterior distribution for $\alpha$. In the case of FIREboxHR we do use the binned SMFs/UVLFs, but generate posteriors of $\alpha$ by fitting Schechter functions to these SMFs/UVLFs using the same approach. The details of our methodology, along with illustrative results, are described in Appendix B.
@@
We estimate the posterior distribution of the slope $\alpha$, which follows an approximately Gaussian distribution (see Appendix B), and calculate its mean and standard deviation for the three samples in each simulation. Figures 9 and 10 show the slopes obtained by our fits for the SMFs and UVLFs respectively, compared to the ones from the proto-MW/LG and the universal SMFs/UVLFs at $z=7$. The all-progenitors scenario is shown in the top panel, while the fossil-record scenario is in the bottom panel. The points and the error bars in each figure denote the mean and standard deviation of the posterior distribution for the low-mass-end/faint-end slope $\alpha$ for the proto-MW/LG (black), Sample I (red), Sample II (blue) and Sample III (green) for each simulation. The mean slope of the universal SMF/UVLF is shown by the yellow line, while the shaded region shows the standard deviation. To estimate population-level parameters, we bootstrap the individual $\alpha$ values, resampling from their Gaussian posteriors. This yields: (1) a central slope estimate ($\tilde{\mu}_{\alpha}$) - median of the bootstrapped means; (2) halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$) - median of the bootstrapped standard deviation. We show these by the horizontal lines and the shaded regions, respectively, in Figures 9 and 10. To compare the results between proto-MW and proto-LG, we separately do the analysis for the isolated and paired galaxy suites. For clarity, we also show these estimates in the rightmost points of both the figures. We emphasize that $\tilde{\sigma}_{\alpha}$ is not the standard deviation of the bootstrapped mean $\alpha$ distribution, but rather an average range of possible $\alpha$ values across all the halos in the respective galaxy suite, defining the halo-to-halo scatter. We discuss the details of our bootstrapping method in Appendix C. Tables 2 and 3 summarizes our slope estimates. We note that the median values for the isolated and paired galaxy suites calculated separately, in general, provide similar conclusions as the median values for the 10 halos combined. However, the results for the paired galaxy suite are derived from only three simulations and thus may not be robust. Since the trend is similar we discuss our results for $\tilde{\mu}_{\alpha}$ and $\tilde{\sigma}_{\alpha}$ when all the halos are combined, unless stated otherwise.
@@
The low-mass-end slope values of the proto-MW/LG and Sample I reported here are different than the ones in Gandhi et al. (2024) by $\approx$0.1, likely due to the difference in the slope estimation method. In addition, some of the present-day intact galaxies in Gandhi et al. (2024) are instead identified as part of our set of disrupted galaxies, so our Sample I is not identical to the set used in that paper.
@@
We find several key features when we include the disrupted galaxy progenitors while computing the slopes at the low-mass/faint end. First, when we compare the individual low-mass-end/faint-end slopes from the SMFs/UVLFs of the three samples with the one from the proto-MW/LG, we find that in most halos the slope of the proto-MW/LG SMF/UVLF is consistently better reproduced by Sample III, followed by Samples II and I, respectively. This is also reflected in the median values, $\tilde{\mu}_{\alpha}$, from our bootstrapped samples. Second, including more disrupted galaxies reduces the standard deviation of $\alpha$ values from Sample I within each halo, by up to $\sim$30% in Sample II and $\sim$50% in Sample III. This is because the increased number of progenitors improves statistical robustness. Third, there is a scatter in the values of the slope across different halos, which implies that different accretion histories can lead to different values of the reconstructed low-mass-end/faint-end slope, leading to poorly constrained values across the realizations. However, we find that this halo-to-halo scatter can be reduced by adding the disrupted galaxies, as shown by the decrease in the median $\tilde{\sigma}_{\alpha}$ values (rightmost points). This reduction in the scatter, in comparison to Sample I, can be $\sim$20% in case of Sample II and $\sim$40% in case of Sample III. The effect of an unusual accretion history in producing a poorly constrained slope can thus be mitigated by adding the disrupted galaxies.
@@
We reach similar conclusions when we consider the effect of adding disrupted galaxies separately to a MW-like environment and a LG-like environment. The halo-to-halo scatter in case of a LG-like environment is smaller compared to that of a MW-like environment, but the results are derived from only three simulations and may not be robust. Additionally, the low-mass-end for Samples I and III is steeper than proto-LG because of the R&R paired halo (see Figure 9). Furthermore, the slopes are susceptible to how present-day and high-$z$ galaxies are selected, which we discuss in VI.2 and hence the should be treated with caution.
@@
Interestingly, the low-mass-end slopes of the SMFs (see Figure 9) and the faint-end slopes of the UVLFs (see Figure 10) differ. The UVLFs have a shallower slope at the faint-end ($\approx$ -1.4) compared to the low-mass-end of the SMFs ($\approx$ -1.7). Even though extrapolations of observed UVLFs predict a faint-end slope of $\approx$ -2, results from FIREboxHR also seem to be pointing to a shallower slope (see Table 2). Although this could depend on the specific distribution used to draw the progenitor masses and the slope estimation method, the difference between the SMF and UVLF slopes invites further investigation. Using the power-law nature of the SMF at the low-mass end and the mass-to-light ratios computed from our progenitor stellar halos, we find that the shallower slope in the UVLFs, can be partially explained due to the superlinear dependence of UV luminosity with stellar mass. Since young, massive O/B stars generate most of the UV radiation and lower-mass galaxies are less efficient at producing them, this leads to suppressed UVLFs at the faint end. We refer the reader to Appendix D for further details and figures demonstrating our analysis. In any case, we better reproduce the proto-MW/LG SMFs/UVLFs after including present-day disrupted galaxies.
@@
Another important difference can be discerned between the all-progenitors scenario and the fossil-record scenario: the latter case results in flatter slopes - by $\sim$0.07 on average across the three samples when both the isolated and paired galaxy suites are combined. As discussed in Section V.3, this effect is due to combining multiple progenitors into a single one in the fossil-record scenario and it gets amplified for Samples II and III as they have more massive substructures with multiple progenitors. The level of flattening varies depending on how both the present-day and high-$z$ progenitor galaxies are selected.
@@
Sample SMFs UVLFs
@@
Sample SMFs UVLFs